OpenStudy (anonymous):

8 3/4 divided by 2 1/6 I got 3 18/26, correct?

OpenStudy (anonymous):

\[8 \frac{ 3 }{ 4 } \div 2 \frac{ 1 }{ 6 } = 3 \frac{ 18 }{ 26 } ?\]

jabez177 (jabez177):

Okay, first of all, 8 x 4 + 3 = what?

OpenStudy (anonymous):

35

OpenStudy (danisaur):

4 1/26

OpenStudy (anonymous):

oh I made a mistake in the problem :o

jabez177 (jabez177):

Okay, so the first one is 35/4. Now, what is 2 x 6 + 1?

OpenStudy (anonymous):

13

OpenStudy (anonymous):

Now I got 4 1/26 I did the problem again

jabez177 (jabez177):

Now that is 13/6 so \[\frac{ 35 }{ 4 }\times \frac{ 6 }{ 13 }\]

jabez177 (jabez177):

Because when you divide fractions, it's the same as Multiplying the Reciprocal.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thank you

jabez177 (jabez177):

What is 35 x 6?